OpenStudy (anonymous):

Find the exact circumference of a circle with the given radius. x/2 miles C =

terenzreignz (terenzreignz):

To get the circumference given the radius, just multiply the radius by \(2\pi\)

Directrix (directrix):

C = 2 * pi * r C = 2* pi * (x/2) C = ? @whizkidzack